OpenStudy (anonymous):

How to factorise the following equation: 2x^2 + 5x - 168

OpenStudy (anonymous):

@Chad123 Do u know discriminant method?

OpenStudy (anonymous):

Could you please show me?

OpenStudy (anonymous):

for quadratic equation Ax^2 + Bx+ C Discriminant, D = sqrt( B^2- 4*A*C) then roots x1 = (- B + D)/ 2*A and x2 = (-B - D)/2*A

OpenStudy (anonymous):

in the problem it is given 2x^2+5x - 168 So A = 2 , B = 5 and C = -168 first find the discriminant and then roots

OpenStudy (anonymous):

@Chad123 can u do it now?

OpenStudy (anonymous):

What is the capital D?

OpenStudy (anonymous):

D is notation for discriminant

OpenStudy (anonymous):

I'll let you know in a second

OpenStudy (anonymous):

The discrimant equalled square root of 1033

OpenStudy (anonymous):

Could you please show me the steps if you don't mind

OpenStudy (anonymous):

u r getting wrong u show ur steps and i will help u in correcting those steps

OpenStudy (anonymous):

Square root 5^2 - 4(2)(-168)

OpenStudy (anonymous):

Is that how you do it?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

what is your second step?

OpenStudy (anonymous):

Adding 25 with 1008

OpenStudy (anonymous):

You get square root of 1033

OpenStudy (anonymous):

4*2*168 = 1344 add 25 with 1344

OpenStudy (anonymous):

No it's equal to 1033

OpenStudy (anonymous):

Sorry your right

OpenStudy (anonymous):

so u got sqrt(1369) which is 37

OpenStudy (anonymous):

now can u find the roots?

OpenStudy (anonymous):

And then I plug it into x1 and x2?

OpenStudy (anonymous):

yah u will get your required roots

OpenStudy (anonymous):

You rock man I'm about to give you a medal

OpenStudy (anonymous):

Thanks for all the help

OpenStudy (anonymous):

:-P

OpenStudy (anonymous):

Ciao :)